# SpecDiff: Accelerating Diffusion Model Inference with Self-Speculation

### v1

@@ Introduction
Towards the advancement of multimodal artificial intelligence, the diffusion model is a typical neural network, achieving remarkable success across various domains (e.g., text-to-image [1] and text-to-video [14] generation), and significantly enabling the rapid development of numerous downstream tasks (e.g., content creation [19]). The inference of the diffusion model is a process of continuously denoising images by iteratively executing the complete model. Each execution is a modification and refinement of the image. Driven by the scaling law [21, 20], the diffusion model with an increasing number of parameters has demonstrated outstanding performance in many scenarios. However, this further incurs significant memory requirements and inference latency. Moreover, the input matrix to the diffusion model inference is transformed from the noise image, which has comparable row values to the model weight, leading to the compute-bound matrix multiplication operator. Reducing the actual computation workload is the direct and effective optimization method for inference acceleration.
@@
Consequently, many previous works have explored techniques to reduce the computation workload, including algorithm optimization, system enhancements, and hardware advancements [2, 3, 4]. Recently, feature caching has become an emerging and promising method that exploits the similarities between features inherent in iterative inference of the diffusion model [16, 17]. By caching and reusing relatively invariant features to replace redundant model computations, it effectively reduces the overall computational workload and achieves significant acceleration. Furthermore, it is evident that the effectiveness of feature caching greatly depends on its design and implementation, specifically, on how to select the features that need to be cached precisely. This selection has a critical impact on both the accuracy and speedup of the diffusion model.
In this paper, we analyze feature caching from a novel perspective, information utilization, and point out that the information utilization is directly related to the design of the feature caching and significantly impacts the overall performance. Current work only focuses on exploiting historical information(e.g., feature similarity between historical steps [16]) for feature caching and fails to predict the future characteristics of the model accurately. As shown in Figure 2(b), historical information only captures the local variation and ignores potential image mutation, resulting in poor performance (over $60\%$ accuracy loss [15]) in some scenarios.
Existing work [29] has proven that the features focused by the diffusion model are strongly related to the time step parameter. This inspires the potential of models with future time steps to focus on information to assist in token selection. From Figure 2(b), the information of the same time step in different iterations is highly similar. Thus we propose a novel paradigm using few self-speculation steps of the original model in advance to introduce future information to capture the global variation. To fully leverage speculative future information for precise feature caching, we present SpecDiff, a training-free multi-level feature caching strategy based on speculative information, including a feature selection algorithm based on self-speculative information and a multi-level feature classification algorithm based on feature importance scores. The contributions are summarized as follows.
@@
  • Extensive experiments show that SpecDiff achieves average $2.80\times$, $2.37\times$ and $3.17\times$ speedup with negligible quality loss on Stable Diffusion 3, 3.5 [22, 23] and FLUX [24] shown in Figure 2(d). As shown in Figure 1, by merging speculative and historical information, SpecDiff overcomes the speedup-accuracy trade-off bottleneck, successfully pushing the Pareto frontier in the efficient diffusion model inference.
@@
As the typical neural network in multimodal artificial intelligence, the diffusion model has seen widespread application across diverse scenarios and has achieved impressive performance. Currently, diffusion transformer (DiT) architectures, represented by Stable Diffusion 3, 3.5 [22, 23], and FLUX [24], have become the mainstream design paradigm for diffusion models. Different from the traditional UNet-based diffusion models, the main composition of the DiT architecture is the transformer block with self-attention mechanism and feed-forward network(FFN) illustrated in Figure 2(a). The self-attention mechanism elevates generation quality via long-range dependency modeling, while the FFN aims to capture inherent features, which compensate for the limitations of the self-attention mechanism. Moreover, the variational auto encoder (VAE) [6] consists of an encoder and a decoder. The encoder is responsible for mapping the image from the pixel space to the latent space, while the decoder performs the reverse mapping. The text encoder and image embedding modules transform the text information and image information into high-dimensional features.
@@
In this paper, we focus on the DiT model inference acceleration. For the general matrix multiplication (GEMM) operator in FFN and self-attention operations of the DiT model, the GEMM ($M\times N\times K$) computation typically involves multiplying an input matrix ($M\times N$) with a weight matrix ($N\times K$). The ratio of computing workload and memory access (also known as arithmetic intensity) is defined as $\frac{M\times N\times K}{M\times K+N\times K}$. As illustrated in Figure 2(a), the input of the DiT backbone in each iterative step is the matrix transformed by the noise image. The rows of the input matrix ($M$) commonly have comparable values with the rows of the weight ($N$) (e.g., $M=4096$ and $N=1536$ in Stable Diffusion 3 [22]), leading to the high arithmetic intensity. The arithmetic intensity reflects the computational efficiency, with higher values indicating that the GEMM is compute-bound rather than memory-bound.
Therefore, most of the works for DiT model acceleration mainly focus on reducing the computation. Model compression (e.g., distillation [13], quantization [12, 25]) exploits the inherent parameter features to make the model smaller, achieving computation reduction, but bringing expensive GPU training hours. Dynamic inference [8, 7] adaptively adjusts the number of model iterations based on the stability of intermediate results over iterations, leading to negligible quality loss but normal acceleration due to the overhead introduction. Feature caching methods [15, 18, 16, 17] cache and reuse the features during iterations based on the similarity between token features, significantly reducing the computation and achieving high speedup with remarkable generation quality.
@@
In this paper, we focus on the optimization of feature caching of the DiT model. Figure 3 shows different designs of feature caching in existing mainstream works. FORA [15] directly caches and reuses the features in adjacent steps, resulting in profound quality loss in most scenarios (e.g., $60\%$ accuracy drop in FID metric). ToCa [18] selects partial important tokens with high attention scores and caches the features of the left tokens in the current step for computation and reuse in the next step. RAS [16] further proposes to utilize the feature similarity from two adjacent steps to achieve the important token selection. By identifying the importance of tokens, these two works have demonstrated notable results in both efficiency and accuracy. Based on the differential similarity of all historical features, TaylorSeer [17] directly approximates the feature in the current step by using features from previous steps.
@@
Inspired by the speculation in LLM [10, 11, 7], we point out that speculation is an effective way to obtain future information by approximating the final result through limited model computation on the same input. Furthermore, as illustrated in Figure 4(a), our method assessing token significance through attention scores reveals significant differences in token importance at time step 2 between consecutive iterations. Historical information can only predict 50% of important tokens. As shown in Figure 4(a), by analyzing high-attention tokens occurring at identical time steps across iterations versus adjacent time steps within the same iteration using a fixed selection proportion, we observe that important tokens cannot be predicted by historical information have a considerable proportion in important tokens predicted by speculative information compared with random prediction. This indicates that tokens deemed significant but unobtainable from historical information can instead be acquired through future information access.
@@
For different cached token ratios CRCR ($0<CR<1$) , we select tokens with the largest importance scores in the quantity of 1CR1-CR as those that need to participate in network computation. For various caching ratios, we find that the number of times tokens are selected presents a significantly skewed distribution, as shown in Figure 5(a). For the case where only 20% of important tokens are selected, the tokens with selection frequencies in the top 25% account for more than 75% of the selections. Moreover, this distribution exhibits a significant long-tail effect: approximately 40% of tokens have never been considered important in dozens of iterations! Tokens that are cached too many times may have tremendous cumulative errors[18]. Therefore, to ensure that these tokens still have the possibility of being selected, we introduce a starvation score into the token’s importance score, which is positively correlated with the number of times the token is cached, denoted as:
@@
After selecting a subset of tokens with the highest scores, how to efficiently approximate the features of the remaining tokens becomes a new key issue. Existing methods adopt the strategy of reusing the relevant strategy from the previous iteration[18, 16], and the following simple equation can describe this process:
@@
Models and Datasets. We evaluated the effectiveness of our method on the advanced DiT text-to-image models, Stable Diffusion 3 [22], Stable Diffusion 3.5 [23], and FLUX.1 Dev [24]. We randomly select 5000 text-image pairs from the COCO 2014 val dataset [5] to generate $\times1024$ images for the experiments.
Metrics and Baseline. To evaluate the quality of the generated images and the alignment between text and images, we chose FID [26], Clip Score [27], and VQA Score [28] as our main experiment evaluation metrics. And for baseline comparison, we compared our method with the best similar method RAS [16], and the SOTA method of feature caching method TaylorSeer [17]. In order to comprehensively compare with the SOTA method TaylorSeer. We add SSIM, PSNR, and memory as new metrics. These baseline methods all have good generation quality with a good acceleration effect. We compared the generation quality of our method with these methods under different acceleration ratios. We set the parameters of baseline models Stable Diffusion 3 [22], 3.5 [23], and FLUX [24] for the default value. The inference step of our experiment is all set to 28. The CFG of Stable Diffusion 3 and 3.5 is set to 7.0, while the FLUX is 3.5.
@@
Speculation Steps Cached Ratio FID $\downarrow$ Clip Score $\uparrow$ VQA Score $\uparrow$ Speedup

### v2

@@ Introduction
Towards the advancement of multimodal artificial intelligence, the diffusion model is a typical neural network, achieving remarkable success across various domains (e.g., text-to-image (zhang2023text) and text-to-video (xing2024survey) generation), and significantly enabling the rapid development of numerous downstream tasks (e.g., content creation (wang2024diffusion)). The inference of the diffusion model is a process of continuously denoising images by iteratively executing the complete model. Each execution is a modification and refinement of the image. Driven by the scaling law (henighan2020scaling; liang2024scaling), the diffusion model with an increasing number of parameters has demonstrated outstanding performance in many scenarios. However, this further incurs significant memory requirements and inference latency. Moreover, the input matrix to the diffusion model inference is transformed from the noise image, which has comparable row values to the model weight, leading to the compute-bound matrix multiplication operator. Reducing the actual computation workload is the direct and effective optimization method for inference acceleration.
@@
Consequently, many previous works have explored techniques to reduce the computation workload, including algorithm optimization, system enhancements, and hardware advancements (ding2025vida; yuan2025vgdfr; zhang2025ditfastattnv2). Recently, feature caching has become an emerging and promising method that exploits the similarities between features inherent in iterative inference of the diffusion model (ras; taylorseer). By caching and reusing relatively invariant features to replace redundant model computations, it effectively reduces the overall computational workload and achieves significant acceleration. Furthermore, it is evident that the effectiveness of feature caching greatly depends on its design and implementation, specifically, on how to select the features that need to be cached precisely. This selection has a critical impact on both the accuracy and speedup of the diffusion model.
In this paper, we analyze feature caching from a novel perspective, information utilization, and point out that the information utilization is directly related to the design of the feature caching and significantly impacts the overall performance. Current work only focuses on exploiting historical information(e.g., feature similarity between historical steps (ras)) for feature caching and fails to predict the future characteristics of the model accurately. As shown in Figure 2(b), historical information only captures the local variation and ignores potential image mutation, resulting in poor performance (over $60\%$ accuracy loss (selvaraju2024fora)) in some scenarios.
Existing work (gao2024mdtv2maskeddiffusiontransformer) has proven that the features focused by the diffusion model are strongly related to the time step parameter. This inspires the potential of models with future time steps to focus on information to assist in token selection. From Figure 2(b), the information of the same time step in different iterations is highly similar. Thus we propose a novel paradigm using few self-speculation steps of the original model in advance to introduce future information to capture the global variation. To fully leverage speculative future information for precise feature caching, we present SpecDiff, a training-free multi-level feature caching strategy based on speculative information, including a feature selection algorithm based on self-speculative information and a multi-level feature classification algorithm based on feature importance scores. The contributions are summarized as follows.
@@
  • Extensive experiments show that SpecDiff achieves average $2.80\times$, $2.37\times$ and $3.17\times$ speedup with negligible quality loss on Stable Diffusion 3, 3.5 (stablediff_3; stable_diff_3_5) and FLUX (FLUX) shown in Figure 2(d). As shown in Figure 1, by merging speculative and historical information, SpecDiff overcomes the speedup-accuracy trade-off bottleneck, successfully pushing the Pareto frontier in the efficient diffusion model inference.
@@
As the typical neural network in multimodal artificial intelligence, the diffusion model has seen widespread application across diverse scenarios and has achieved impressive performance. Currently, diffusion transformer (DiT) architectures, represented by Stable Diffusion 3, 3.5 (stablediff_3; stable_diff_3_5), and FLUX (FLUX), have become the mainstream design paradigm for diffusion models. Different from the traditional UNet-based diffusion models, the main composition of the DiT architecture is the transformer block with self-attention mechanism and feed-forward network(FFN) illustrated in Figure 2(a). The self-attention mechanism elevates generation quality via long-range dependency modeling, while the FFN aims to capture inherent features, which compensate for the limitations of the self-attention mechanism. Moreover, the variational auto encoder (VAE) (kingma2013auto) consists of an encoder and a decoder. The encoder is responsible for mapping the image from the pixel space to the latent space, while the decoder performs the reverse mapping. The text encoder and image embedding modules transform the text information and image information into high-dimensional features.
@@
In this paper, we focus on the DiT model inference acceleration. For the general matrix multiplication (GEMM) operator in FFN and self-attention operations of the DiT model, the GEMM ($M\times N\times K$) computation typically involves multiplying an input matrix ($M\times N$) with a weight matrix ($N\times K$). The ratio of computing workload and memory access (also known as arithmetic intensity) is defined as $\frac{M\times N\times K}{M\times K+N\times K}$. As illustrated in Figure 2(a), the input of the DiT backbone in each iterative step is the matrix transformed by the noise image. The rows of the input matrix ($M$) commonly have comparable values with the rows of the weight ($N$) (e.g., $M=4096$ and $N=1536$ in Stable Diffusion 3 (stablediff_3)), leading to the high arithmetic intensity. The arithmetic intensity reflects the computational efficiency, with higher values indicating that the GEMM is compute-bound rather than memory-bound.
Therefore, most of the works for DiT model acceleration mainly focus on reducing the computation. Model compression (e.g., distillation (yin2024one), quantization (wu2025quantcache; wang2025maketrainingflexibledeploymentefficient)) exploits the inherent parameter features to make the model smaller, achieving computation reduction, but bringing expensive GPU training hours. Dynamic inference (zhang2023adadiff; xu2025specee) adaptively adjusts the number of model iterations based on the stability of intermediate results over iterations, leading to negligible quality loss but normal acceleration due to the overhead introduction. Feature caching methods (selvaraju2024fora; toca; ras; taylorseer) cache and reuse the features during iterations based on the similarity between token features, significantly reducing the computation and achieving high speedup with remarkable generation quality.
@@
In this paper, we focus on the optimization of feature caching of the DiT model. Figure 3 shows different designs of feature caching in existing mainstream works. FORA (selvaraju2024fora) directly caches and reuses the features in adjacent steps, resulting in profound quality loss in most scenarios (e.g., $60\%$ accuracy drop in FID metric). ToCa (toca) selects partial important tokens with high attention scores and caches the features of the left tokens in the current step for computation and reuse in the next step. RAS (ras) further proposes to utilize the feature similarity from two adjacent steps to achieve the important token selection. By identifying the importance of tokens, these two works have demonstrated notable results in both efficiency and accuracy. Based on the differential similarity of all historical features, TaylorSeer (taylorseer) directly approximates the feature in the current step by using features from previous steps.
@@
Inspired by the speculation in LLM (li2024eagle; li2024specpim; xu2025specee), we point out that speculation is an effective way to obtain future information by approximating the final result through limited model computation on the same input. Furthermore, as illustrated in Figure 4(a), our method assessing token significance through attention scores reveals significant differences in token importance at time step 2 between consecutive iterations. Historical information can only predict 50% of important tokens. As shown in Figure 4(a), by analyzing high-attention tokens occurring at identical time steps across iterations versus adjacent time steps within the same iteration using a fixed selection proportion, we observe that important tokens cannot be predicted by historical information have a considerable proportion in important tokens predicted by speculative information compared with random prediction. This indicates that tokens deemed significant but unobtainable from historical information can instead be acquired through future information access.
@@
For different cached token ratios CRCR ($0<CR<1$) , we select tokens with the largest importance scores in the quantity of 1CR1-CR as those that need to participate in network computation. For various caching ratios, we find that the number of times tokens are selected presents a significantly skewed distribution, as shown in Figure 5(a). For the case where only 20% of important tokens are selected, the tokens with selection frequencies in the top 25% account for more than 75% of the selections. Moreover, this distribution exhibits a significant long-tail effect: approximately 40% of tokens have never been considered important in dozens of iterations! Tokens that are cached too many times may have tremendous cumulative errors(toca). Therefore, to ensure that these tokens still have the possibility of being selected, we introduce a starvation score into the token’s importance score, which is positively correlated with the number of times the token is cached, denoted as:
@@
After selecting a subset of tokens with the highest scores, how to efficiently approximate the features of the remaining tokens becomes a new key issue. Existing methods adopt the strategy of reusing the relevant strategy from the previous iteration(toca; ras), and the following simple equation can describe this process:
@@
Models and Datasets. We evaluated the effectiveness of our method on the advanced DiT text-to-image models, Stable Diffusion 3 (stablediff_3), Stable Diffusion 3.5 (stable_diff_3_5), and FLUX.1 Dev (FLUX). We randomly select 5000 text-image pairs from the COCO 2014 val dataset (lin2014microsoft) to generate $\times1024$ images for the experiments.
Metrics and Baseline. To evaluate the quality of the generated images and the alignment between text and images, we chose FID (DBLP:journals/corr/HeuselRUNKH17), Clip Score (radford2021learningtransferablevisualmodels), and VQA Score (lin2024evaluatingtexttovisualgenerationimagetotext) as our main experiment evaluation metrics. And for baseline comparison, we compared our method with the best similar method RAS (ras), and the SOTA method of feature caching method TaylorSeer (taylorseer). In order to comprehensively compare with the SOTA method TaylorSeer. We add SSIM, PSNR, and memory as new metrics. These baseline methods all have good generation quality with a good acceleration effect. We compared the generation quality of our method with these methods under different acceleration ratios. We set the parameters of baseline models Stable Diffusion 3 (stablediff_3), 3.5 (stable_diff_3_5), and FLUX (FLUX) for the default value. The inference step of our experiment is all set to 28. The CFG of Stable Diffusion 3 and 3.5 is set to 7.0, while the FLUX is 3.5.
@@
Spec. Steps Cached Ratio FID $\downarrow$ Clip Score $\uparrow$ VQA Score $\uparrow$ Speedup
@@ Conclusion
Acknowledgments
This work was sponsored by Shanghai Rising-Star Program (No. 24QB2706200) and the National Natural Science Foundation of China (No. U21B2031).